What are the empirical and molecular formulas of a compound that is 40.27% potassium, 26.78% chromium, and 32.95% oxygen? The molar mass of the compound is 194.2 g/mol.

1
Expert's answer
2022-02-14T22:04:01-0500

The mass of each element is as shown below;

mass of K = 40.27g

mass of Cr = 26.78g

mass of O = 32.95g

Find the moles of each element

moles of K = "40.27\/39.10=1.030"

moles of Cr = "26.78\/52.00=0.5150"

moles of O = "32.95\/16.00=2.059"

Let the empirical formula be "K_xCr_yO_z"

divide the number of moles of each element by the least number of moles.

"x=1.030\/0.5150=2"

"y=0.5150\/0.5150=1"

"z=2.059\/0.5150=4"

Hence the empirical formula is "K_2CrO_4"

molecular formula "=(K_2CrO_4)n"

empirical mass ="39.1\times 2 +52\times 1 +16\times 4=194.2g/mol"

"194.2=194.2n"

"n=1"

Therefore, the molecular formula ="K_2CrO_4"